# FantasyStyle: Controllable Stylized Distillation for 3D Gaussian Splatting

### v1

@@ Abstract
The success of 3DGS in generative and editing applications has sparked growing interest in 3DGS-based style transfer. However, current methods still face two major challenges: (1) multi-view inconsistency often leads to style conflicts, resulting in appearance smoothing and distortion; and (2) heavy reliance on VGG features, which struggle to disentangle style and content from style images, often causing content leakage and excessive stylization. To tackle these issues, we introduce FantasyStyle, a 3DGS-based style transfer framework, and the first to rely entirely on diffusion model distillation. It comprises two key components: (1) Multi-View Frequency Consistency. We enhance cross-view consistency by applying a 3D filter to multi-view noisy latent, selectively reducing low-frequency components to mitigate stylized prior conflicts. (2) Controllable Stylized Distillation. To suppress content leakage from style images, we introduce negative guidance to exclude undesired content. In addition, we identify the limitations of Score Distillation Sampling and Delta Denoising Score in 3D style transfer and remove the reconstruction term accordingly. Building on these insights, we propose a controllable stylized distillation that leverages negative guidance to more effectively optimize the 3D Gaussians. Extensive experiments demonstrate that our method consistently outperforms state-of-the-art approaches, achieving higher stylization quality and visual realism across various scenes and styles.
@@ 1 Introduction
With the widespread adoption of 3D content in virtual reality (VR), augmented reality (AR), and related fields, user expectations for visual quality have shifted from traditional photorealism toward more artistic and personalized styles. In this context, style transfer has emerged as a key technique for altering the visual appearance of 3D content, attracting increasing attention from the research community. Previous works (Zhang et al. 2022; Jung et al. 2024; Zhang, Fernandez-Labrador, and Schroers 2024; Zhang et al. 2023a) have explored 3D style transfer based on Neural Radiance Fields (NeRF) (Mildenhall et al. 2021), typically by introducing VGG-based stylization losses to ensure both style consistency and structural coherence in rendered views. However, NeRF suffers from slow rendering and long training times, limiting its practical applications. In contrast, 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023), as an emerging 3D representation, has become a research focus for 3D style transfer due to its significantly faster rendering and superior visual quality.
@@
Most recent studies on style transfer for 3DGS are based on VGG features, drawing inspiration from the early work on feature statistic matching (Gatys, Ecker, and Bethge 2016). These methods (Zhang et al. 2024; Saroha et al. 2024a; Kovács, Hermosilla, and Raidou 2024) typically extract VGG features and compute Gram matrices to achieve texture style transfer. Meanwhile, some studies (Zuo et al. 2024; Gu et al. 2024) attempt to leverage diffusion models to generate stylized 2D images, which are then further optimized back into the 3D scene to enable style transfer. Although these methods have shown promising results, they still face two critical challenges. (1) Multi-view inconsistency. Using 2D diffusion priors to guide 3D stylization often leads to inconsistent appearances across views, especially between adjacent ones. These inconsistencies can lead to style conflicts, hinder optimization, and distort geometry and appearance, as shown in Fig. 4 (b). (2) Content leakage from style images and over-stylization. As shown in Fig. 1 (A) (B), VGG-based methods struggle to separate content and style, often causing unintended content transfer from the style image. Their emphasis on low-level textures often results in over-stylization, which obscures structural details.
@@
To address the aforementioned challenges, we propose FantasyStyle, a 3DGS-based style transfer framework. Drawing inspiration from FreeU (Si et al. 2024) and FreeInit (Wu et al. 2024), which emphasize the critical role of frequency components in image and video generation, we perform a frequency-domain analysis of 2D multi-view stylized priors. We observe that low-frequency components primarily reflect view-dependent local details and tend to exhibit poor consistency across different viewpoints. In contrast, high-frequency components more stably capture texture features and demonstrate better consistency across views. Therefore, appropriately attenuating the low-frequency components can help alleviate inter-view discrepancies and improve overall multi-view consistency. Based on this observation, we introduce Multi-View Frequency Consistency (MVFC), which applies a 3D frequency-domain filter to the multi-view latent after DDIM noise injection. MVFC preserves all high-frequency components and selectively suppresses low-frequency components, effectively reducing stylized prior conflicts and improving both consistency and stylization quality.
@@
In parallel, to eliminate potential content leakage from the style image, we incorporate negative guidance during the denoising phase, ensuring that the derived 2D stylized prior remains free of content information. While methods such as Score Distillation Sampling (SDS) and Delta Denoising Score (DDS) have been employed for 3D scene optimization, they are not directly applicable in our context, as they tend to produce overly smooth images that lose important brushstroke details present in style images. To overcome this limitation, we remove the reconstruction term from these objectives and incorporate negative guidance to design a Controllable Stylized Distillation (CSD), which leverages 2D stylized priors to facilitate faithful and structurally coherent style transfer in 3D scenes.
@@ 2 Related Work
Figure 2: Overview of our proposed method. Inspired by DDS (Hertz, Aberman, and Cohen-Or 2023), FantasyStyle consists of two distinct pathways: Source Image and Rendered Image, each highlighted in different colors. We propose Controllable Stylized Distillation (CSD) to optimize the 3D scene. In Rendered Image pathway, we introduce a multi-view frequency consistency (MVFC) and inject style image features to obtain a multi-view consistent 2D stylized prior. Additionally, we incorporate negative guidance to suppress potential content leakage from the style image.
@@
2D Style Transfer Based on Diffusion Models. Style transfer aims to separate style and content representations from a style image and apply the style to another image. With the rise of diffusion models, 2D style transfer has seen notable advances. Recent studies leverage cross-attention (Zhang et al. 2023b) and self-attention mechanisms (Hertz et al. 2024; Chung, Hyun, and Heo 2024) to inject style features into the generation process. In addition, frequency-domain techniques (Xu et al. 2025; Gao et al. 2024) have been incorporated into the diffusion process, effectively enhancing the model’s ability to control stylistic attributes. More recently, methods (Frenkel et al. 2024; Ouyang, Li, and Hou 2025) based on Low-Rank Adaptation (LoRA) (Hu et al. 2022) have been proposed to achieve a more effective separation of content and style features, thereby improving the flexibility and diversity of style transfer. We propose a Controllable Stylized Distillation approach that utilizes 2D stylized priors to optimize 3D scenes, effectively bridging the gap between 2D and 3D style transfer.
@@
3DGS Style Transfer Based on VGG. 3D style transfer is a conditional generative task that applies the visual style of a 2D image to a 3D representation while preserving the original geometric structure. The VGG network, one of the earliest architectures used in style transfer, remains a widely adopted feature extractor in both 2D and 3D tasks. Saroha et al.(Saroha et al. 2024b) employ VGG and AdaIN to transfer styles to specific views. StyleSplat(Jain et al. 2024) introduces a nearest neighbor feature matching loss between VGG features of rendered and reference images. SGSST (Galerne et al. 2024) proposes a Simultaneously Optimized Scales loss to enable ultra-high-resolution 3D stylization. StyleGaussian (Liu et al. 2024) embeds 2D VGG scene features into transformed 3D Gaussian features, which are fused with reference features and decoded into stylized RGB outputs for more coherent, view-consistent stylization. However, these methods often suffer from content leakage from the style image. To address this, we propose a negative guidance mechanism to suppress the transfer of irrelevant content information from the style image.
@@
3DGS Style Transfer Based on Diffusion Model. With the powerful generative capacity of text-to-image (T2I) diffusion models, recent 3D style transfer methods increasingly incorporate pre-trained diffusion priors to guide stylization. InstantStyleGaussian (Yu et al. 2024) accelerates the process by generating target-style images via diffusion and integrating them into the training set to iteratively optimize the 3D Gaussian Splatting scene. ArtNVG (Gu et al. 2024) enhances content-style disentanglement using CSGO (Xing et al. 2024) and Tile-ControlNet (Zhang, Rao, and Agrawala 2023), and introduces an attention-based view alignment module for consistent local texture and color. StyleMe3D (Zhuang et al. 2025) proposes dynamic style score distillation in the latent space of Stable Diffusion for improved semantic alignment. These methods, when guided by 2D images for 3D scene stylization, do not explicitly enforce multi-view consistency. To address this, we propose a Multi-View Frequency Consistency to improve consistency across views.
@@ 3 Preliminary
Score Distillation Sampling (SDS) (Poole et al. 2022) is widely used in 3D generation, leveraging pretrained diffusion models with rich 2D priors to distill knowledge into 3DGS models. Ignoring the UNet Jacobian, it is formulated as:
@@
Delta Denoising Score (DDS) (Hertz, Aberman, and Cohen-Or 2023) extends SDS for editing tasks by estimating the score difference between target and source distributions to guide updates. This mitigates the noisy gradients of SDS that often lead to blurred or unclear images. It is defined as:
@@ 4.1 Multi-View Frequency Consistency
3D Frequency Analysis. Inspired by FreeU (Si et al. 2024) and FreeInit (Wu et al. 2024), which emphasize the critical role of frequency components in image and video generation, we perform a frequency-domain analysis of the multi-view stylized priors to better understand their behavior. To visualize more intuitively, we use IP-Adapter (Ye et al. 2023) to generate 2D multi-view stylized images. Moreover, since our task only involves fine-tuning color parameters to achieve texture-level style consistency without the need to preserve geometric structure, we introduce ControlNet to help guide the structural information in the images. Given $N$ rendered views from a 3D scene, we first apply DDIM (Ho, Jain, and Abbeel 2020) to add noise and obtain $z_{t}$:
@@ 4.2 Controllable Stylized Distillation
Style Injection. As shown in Fig. 2, we build upon Stable Diffusion XL (SDXL) and incorporate IP-Adapter for style injection, aiming to obtain 2D stylized priors. Although the 3D scene’s geometry remains fixed during optimization, generating 2D stylized priors often causes geometric information loss, leading to structural inconsistencies across views. To address this, we incorporate ControlNet (Zhang, Rao, and Agrawala 2023) to explicitly guide structural features during stylization, improving the geometric consistency of the 2D stylized priors.
@@
where $\mathcal{P}$ denotes the text prompt, $\varnothing$ represents the null-text embedding, and $\beta$ is the CFG scale. To suppress specific content information, $\varnothing$ is replaced with the negative prompt $\mathcal{P}_{neg}$, thereby enabling negative guidance. Negative prompts can be generated using large multimodal language models. For convenience, we leverage the IPAdapter-Instruct (IP) (Rowles et al. 2024) to extract both style and content features from the style image, which are then used to construct the negative prompt. Accordingly, Eq. 8 is reformulated as follows:
@@
$\beta\left(\epsilon_{\phi}(z_{t},t,[\mathcal{P},\text{IP}(I_{r})^{s}])-\epsilon_{\theta}(z_{t},t,\text{IP}(I_{r})^{c})\right),$
@@ 5.1 Experimental Setup
Datasets. Our evaluation utilizes several distinct scene datasets (IN2N (Haque et al. 2023), Tandt DB (Kerbl et al. 2023), and Mip-NeRF 360 (Barron et al. 2022)), with 8 different image styles applied to each scene.
Metrics. We use Art-FID (Wright and Ommer 2022) to jointly evaluate content and style preservation. LPIPS (Zhang et al. 2018) and RMSE measure short-term and long-term content consistency with respect to the input, while FID assesses both style fidelity to the style image and content preservation relative to the original input.
Comparison Methods. We compare our method with recent state-of-the-art 3DGS-based style transfer methods, including SGSST (Galerne et al. 2024) and StyleGaussian (Liu et al. 2024).
Implementation Detail. We deploy 3DGS, SDXL (Podell et al. 2023), ControlNet (Zhang, Rao, and Agrawala 2023), and IP-Adapter (Ye et al. 2023) in all experiments. DDIM is used for noise sampling, with the CFG scale $\beta$ set to 7.5. For MVFC, the parameter $\gamma$ is set to 0.9. All experiments are conducted on 2 NVIDIA L20 (48GB) GPUs.
@@ 5.4 Limitations
Our method is not real-time and requires retraining for each new style. While optimization based on SDXL can be time-consuming, the cost can be mitigated by employing smaller diffusion models, reducing resolution, or tuning the learning rate for color-related parameters.
@@ 6 Conclusion
We propose FantasyStyle, a 3DGS-based style transfer framework that fully leverages diffusion model distillation. It consists of two key components: (1) Multi-View Frequency Consistency, which applies a 3D frequency-domain filter to DDIM-noised multi-view latent to preserve high-frequency details and selectively suppress low-frequency for improved view consistency; and (2) Controllable Stylized Distillation, which introduces negative guidance to prevent content leakage and leverages 2D stylized priors to optimize the 3D scene, enabling effective 3D style transfer. Unlike VGG-based methods, FantasyStyle bridges 2D and 3D stylization, enabling flexible adaptation of 2D techniques to 3D scenes. Experiments show that our method outperforms existing approaches in visual quality and content preservation.

### v2

@@ Abstract
The success of 3DGS in generative and editing applications has sparked growing interest in 3DGS-based style transfer. However, current methods still face two major challenges: (1) multi-view inconsistency often leads to style conflicts, resulting in appearance smoothing and distortion; and (2) heavy reliance on VGG features, which struggle to disentangle style and content from style images, often causing content leakage and excessive stylization. To tackle these issues, we introduce FantasyStyle, a 3DGS-based style transfer framework, and the first to rely entirely on diffusion model distillation. It comprises two key components: (1) Multi-View Frequency Consistency. We enhance cross-view consistency by applying a 3D filter to multi-view noisy latent, selectively reducing low-frequency components to mitigate stylized prior conflicts. (2) Controllable Stylized Distillation. To suppress content leakage from style images, we introduce negative guidance to exclude undesired content. In addition, we identify the limitations of Score Distillation Sampling and Delta Denoising Score in 3D style transfer and remove the reconstruction term accordingly. Building on these insights, we propose a controllable stylized distillation that leverages negative guidance to more effectively optimize the 3D Gaussians. Extensive experiments demonstrate that our method consistently outperforms state-of-the-art approaches, achieving higher stylization quality and visual realism across various scenes and styles. The code is available at https://github.com/yangyt46/FantasyStyle.
@@ 1 Introduction
With the widespread adoption of 3D content (he2025survey; ding2025mevis; wang2025point; wang2025taylor) in virtual reality (VR), augmented reality (AR), and related fields, user expectations for visual quality have shifted from traditional photorealism toward more artistic and personalized styles. In this context, style transfer has emerged as a key technique for altering the visual appearance of 3D content, attracting increasing attention from the research community. Previous works (zhang2022arf; jung2024geometry; zhang2024coarf; zhang2023ref) have explored 3D style transfer based on Neural Radiance Fields (NeRF) (mildenhall2021nerf), typically by introducing VGG-based stylization losses to ensure both style consistency and structural coherence in rendered views. However, NeRF suffers from slow rendering and long training times, limiting its practical applications. In contrast, 3D Gaussian Splatting (3DGS) (kerbl20233d), as an emerging 3D representation, has become a research focus for 3D style transfer due to its faster rendering and superior visual quality.
@@
Most recent studies on style transfer for 3DGS are based on VGG features, drawing inspiration from the early work on feature statistic matching (gatys2016image). These methods (zhang2024stylizedgs; saroha2024gaussian; kovacs2024g) typically extract VGG features and compute Gram matrices to achieve texture style transfer. Meanwhile, some studies (zuo2024towards; ArtNVG) attempt to leverage diffusion models to generate stylized 2D images, which are then further optimized back into the 3D scene to enable style transfer. Although these methods have shown promising results, they still face two critical challenges. (1) Multi-view inconsistency. Using 2D diffusion priors to guide 3D stylization often leads to inconsistent appearances across views, especially between adjacent ones. These inconsistencies can lead to style conflicts, hinder optimization, and distort geometry and appearance, as shown in Fig. 4 (b). (2) Content leakage from style images and over-stylization. As shown in Fig. 1 (A) (B), VGG-based methods struggle to separate content and style, often causing unintended content transfer from the style image. Their emphasis on low-level textures often results in over-stylization, which obscures structural details.
@@
To address the aforementioned challenges, we propose FantasyStyle, a 3DGS-based style transfer framework. Drawing inspiration from FreeU (si2024freeu) and FreeInit (wu2024freeinit), which emphasize the critical role of frequency components in image and video generation, we perform a frequency-domain analysis of 2D multi-view stylized priors. We observe that low-frequency components primarily reflect view-dependent local details and tend to exhibit poor consistency across different viewpoints. In contrast, high-frequency components more stably capture texture features and demonstrate better consistency across views. Therefore, appropriately attenuating the low-frequency components can help alleviate inter-view discrepancies and improve overall multi-view consistency. Based on this observation, we introduce Multi-View Frequency Consistency (MVFC), which applies a 3D frequency-domain filter to the multi-view latent after DDIM noise injection. MVFC preserves all high-frequency components and selectively suppresses low-frequency components, effectively reducing stylized prior conflicts and improving both consistency and stylization quality.
@@
In parallel, to eliminate potential content leakage from the style image, we incorporate negative guidance during the denoising phase, ensuring that the derived 2D stylized prior remains free of content information. While methods (chung2023luciddreamer) such as Score Distillation Sampling (SDS) and Delta Denoising Score (DDS) have been employed for 3D scene optimization, they are not directly applicable in our context, as they tend to produce overly smooth images that lose important brushstroke details present in style images. To overcome this limitation, we remove the reconstruction term from these objectives and incorporate negative guidance to design a Controllable Stylized Distillation (CSD), which leverages 2D stylized priors to facilitate faithful and structurally coherent style transfer in 3D scenes.
@@ 2 Related Work
Figure 2: Overview of our proposed method. Inspired by DDS (hertz2023delta), FantasyStyle consists of two distinct pathways: Source Image and Rendered Image, each highlighted in different colors. We propose Controllable Stylized Distillation (CSD) to optimize the 3D scene. In Rendered Image pathway, we introduce a multi-view frequency consistency (MVFC) and inject style image features to obtain a multi-view consistent 2D stylized prior. Additionally, we incorporate negative guidance to suppress potential content leakage from the style image.
@@
2D Style Transfer Based on Diffusion Models. Style transfer aims to separate style and content representations from a style image and apply the style to another image. With the rise of diffusion models, 2D style transfer has seen notable advances. Recent studies leverage cross-attention (zhang2023inversion) and self-attention mechanisms (hertz2024style; chung2024style) to inject style features into the generation process. In addition, frequency-domain techniques (xu2025stylessp; gao2024frequency) have been incorporated into the diffusion process, effectively enhancing the model’s ability to control stylistic attributes. More recently, methods (frenkel2024implicit; ouyang2025k) based on Low-Rank Adaptation (LoRA) (hu2022lora) have been proposed to achieve a more effective separation of content and style features, thereby improving the flexibility and diversity of style transfer. We propose a Controllable Stylized Distillation approach that utilizes 2D stylized priors to optimize 3D scenes, effectively bridging the gap between 2D and 3D style transfer.
@@
3DGS Style Transfer Based on VGG. 3D style transfer is a conditional generative task that applies the visual style of a 2D image to a 3D representation while preserving the original geometric structure. The VGG network, one of the earliest architectures used in style transfer, remains a widely adopted feature extractor in both 2D and 3D tasks. Saroha et al.(saroha2024gaussiansplattingstyle) employ VGG and AdaIN to transfer styles to specific views. StyleSplat(StyleSplat) introduces a nearest neighbor feature matching loss between VGG features of rendered and reference images. SGSST (galerne2024sgsst) proposes a Simultaneously Optimized Scales loss to enable ultra-high-resolution 3D stylization. StyleGaussian (StyleGaussian) embeds 2D VGG scene features into transformed 3D Gaussian features, which are fused with reference features and decoded into stylized RGB outputs for more coherent, view-consistent stylization. However, these methods often suffer from content leakage from the style image. To address this, we propose a negative guidance mechanism to suppress the transfer of irrelevant content information from the style image.
@@
3DGS Style Transfer Based on Diffusion Model. With the powerful generative capacity of text-to-image (T2I) diffusion models, recent 3D style transfer methods increasingly incorporate pre-trained diffusion priors to guide stylization. InstantStyleGaussian (InstantStyleGaussian) accelerates the process by generating target-style images via diffusion and integrating them into the training set to iteratively optimize the 3D Gaussian Splatting scene. ArtNVG (ArtNVG) enhances content-style disentanglement using CSGO (xing2024csgo) and Tile-ControlNet (zhang2023adding), and introduces an attention-based view alignment module for consistent local texture and color. StyleMe3D (zhuang2025styleme3d) proposes dynamic style score distillation in the latent space of Stable Diffusion for improved semantic alignment. These methods, when guided by 2D images for 3D scene stylization, do not explicitly enforce multi-view consistency. To address this, we propose a Multi-View Frequency Consistency to improve consistency across views.
@@ 3 Preliminary
Score Distillation Sampling (SDS) (poole2022dreamfusion) is widely used in 3D generation, leveraging pretrained diffusion models with rich 2D priors to distill knowledge into 3DGS models. Ignoring the UNet Jacobian, it is formulated as:
@@
Delta Denoising Score (DDS) (hertz2023delta) extends SDS for editing tasks by estimating the score difference between target and source distributions to guide updates. This mitigates the noisy gradients of SDS that often lead to blurred or unclear images. It is defined as:
@@ 4.1 Multi-View Frequency Consistency
3D Frequency Analysis. Inspired by FreeU (si2024freeu) and FreeInit (wu2024freeinit), which emphasize the critical role of frequency components in image and video generation, we perform a frequency-domain analysis of the multi-view stylized priors to better understand their behavior. To visualize more intuitively, we use IP-Adapter (ye2023ip) to generate 2D multi-view stylized images. Moreover, since our task only involves fine-tuning color parameters to achieve texture-level style consistency without the need to preserve geometric structure, we introduce ControlNet to help guide the structural information in the images. Given $N$ rendered views from a 3D scene, we first apply DDIM (ho2020denoising) to add noise and obtain $z_{t}$:
@@ 4.2 Controllable Stylized Distillation
Style Injection. As shown in Fig. 2, we build upon Stable Diffusion XL (SDXL) and incorporate IP-Adapter for style injection, aiming to obtain 2D stylized priors. Although the 3D scene’s geometry remains fixed during optimization, generating 2D stylized priors often causes geometric information loss, leading to structural inconsistencies across views. To address this, we incorporate ControlNet (zhang2023adding) to explicitly guide structural features during stylization, improving the geometric consistency of the 2D stylized priors.
@@
where $\mathcal{P}$ denotes the text prompt, $\varnothing$ represents the null-text embedding, and $\beta$ is the CFG scale. To suppress specific content information, $\varnothing$ is replaced with the negative prompt $\mathcal{P}_{neg}$, thereby enabling negative guidance. Negative prompts can be generated using large multimodal language models. For convenience, we leverage the IPAdapter-Instruct (IP) (rowles2024ipadapter) to extract both style and content features from the style image, which are then used to construct the negative prompt. Accordingly, Eq. 8 is reformulated as follows:
@@
$\beta\left(\epsilon_{\phi}(z_{t},t,[\mathcal{P},\text{IP}(I_{r})^{s}])-\epsilon_{\phi}(z_{t},t,\text{IP}(I_{r})^{c})\right),$
@@ 5.1 Experimental Setup
Datasets. Our evaluation utilizes several distinct scene datasets (IN2N (haque2023instruct), Tandt DB (kerbl20233d), and Mip-NeRF 360 (barron2022mip)), with 8 different image styles applied to each scene.
Metrics. We use Art-FID (wright2022artfid) to jointly evaluate content and style preservation. LPIPS (zhang2018unreasonable) and RMSE measure short-term and long-term content consistency with respect to the input, while FID assesses both style fidelity to the style image and content preservation relative to the original input.
Comparison Methods. We compare our method with recent state-of-the-art 3DGS-based style transfer methods, including SGSST (galerne2024sgsst) and StyleGaussian (StyleGaussian).
Implementation Detail. We deploy 3DGS, SDXL (podell2023sdxl), ControlNet (zhang2023adding), and IP-Adapter (ye2023ip) in all experiments. DDIM is used for noise sampling, with the CFG scale $\beta$ set to 7.5. For MVFC, the parameter $\gamma$ is set to 0.9. All experiments are conducted on 2 NVIDIA L20 (48GB) GPUs.
@@ 5.4 Limitations
Although SDXL-based optimization is time-consuming, it can be mitigated by using smaller models, lower resolutions, or adjusted color-related learning rates.
@@ 6 Conclusion
We propose FantasyStyle, a 3DGS-based style transfer framework that fully exploits diffusion model distillation. It features two key components: (1) Multi-View Frequency Consistency, which applies a 3D frequency filter to DDIM-noised multi-view latents to preserve high-frequency details and suppress low-frequency components for better view consistency; and (2) Controllable Stylized Distillation, which introduces negative guidance to prevent content leakage and uses 2D stylized priors to optimize the 3D scene. Experiments show that our method outperforms existing approaches in visual quality and content preservation.
Acknowledgments
This project is sponsored by Shanghai Pujiang Programme 24PJD030 and Natural Science Foundation of Shanghai 25ZR1402138. Changshuo Wang is supported by the European Union’s Horizon 2024 Research and Innovation Programme for the Marie Skłodowska-Curie Actions under Grant No. 101211118.